# Improving Image Captioning Descriptiveness by Ranking and LLM-based Fusion

### v1

@@ 5 Evaluation
In this section, we investigate whether the best SoTA model, specifically OFA, can learn to generate more descriptive captions when trained using our generated captions as groun-truths. To conduct this experiment, we finetune the model for the image captioning task on MS-COCO by replacing the original ground-truths with the captions generated by our model. Due to limited computational resources (see Appendix A.3 for details), we utilize the Base version of OFA and follow the setup parameters provided by the authors. The 5000 samples from the MS-COCO test set are randomly divided into training (3500 samples), validation (500 samples), and testing (1000 samples). To minimize bias introduced by partitioning, we repeat this split process five times. Table 4 reports the quality captioning performance for OFABase trained on the original MS-COCO and finetuned on MS-COCO with our ground-truths, respectively. It is possible to see that the results for the finetuned model are slightly higher than the baseline ones, i.e. 2 for B@4, 1 for M and 10 for C, respectively.
Table 4: Captioning performance comparison between two versions of OFABase, i.e. the model pretrained on MS-COCO (as known as the baseline) and the model finetuned on the dataset with our generated captions used as ground-truth. It is important to note that the reported performances represent the mean and standard deviation across 5 train-val-test splits.
@@ 6 Discussion
Existing datasets and caption models have several limitations, such as short ground-truth captions and generated captions that tend to look alike and are semantically poor. In this article we propose a caption model that tries to overcome the previous limitations by generating longer and more detailed captions without requiring human intervention. The captions generated by our model were evaluated from different points of view. Here, we summarize the contributions of our work, answering the research questions posed in the introduction. In section 4.1, we address [RQ1] and demonstrate that the captions generated by our proposed model outperform those generated by the SoTA models in terms of accuracy, diversity and semantic richness. In section 4.2, we tackle [RQ2] through a subjective study involving human participants. Based on human judgment, our model generates captions that are more appealing than those produced by SoTA models. In section 5, we aim to answer [RQ3]. The results of our experiment highlight how current image captioning models lean towards the use of common words, phrases and linguistic patterns. Consequently, this leads to a dearth diversity in the generated captions, even when the models are trained on more elaborate and descriptive caption datasets Chen et al. (2022). This phenomenon, known as the mode collapse problem, has been widely discussed in generative modeling Zhou et al. (2020). Addressing this issue requires further investment. Consequently, our proposed caption model can serve as a valuable resource for generating ground-truth captions and pushing models beyond their current limitations.
Societal impact.
@@
Compared to previous image captioning models, our model demonstrates improvements in performance and is more suitable for helping people with visual impairments. The models that are exploited by our proposal are pretrained on large-scale data, and the data are not guaranteed to contain no toxic contents, which could negatively impact the output generated. While we have observed only a small number of such cases in our qualitative analysis, it is imperative to exercise caution when deploying the model into practical applications. Furthermore, more research investigations are needed to ensure better control over the model output.
Limitations.
@@
The limitations of the proposed model are mainly three. The first is that sometimes the five SoTA captioners generate identical or very similar captions (this occurs for about 150 of the 5000 images in the MS-COCO test set). In this case the fusion of the captions collapses into a caption that does not offer more detail. An example is the following where BLIP-2, ExpNet-v2 and OFA generate the same caption, namely “a giraffe standing in the middle of a field”, GIT describes the image content as “a giraffe standing in a field with trees in the background”, finally ViT-GPT2 predicts that “a giraffe standing in a field with trees”. The caption generated by our model is “a giraffe stands majestically in the middle of a field”. Secondly, for 16 test images, the LLM model generates captions with the prefix “The caption for the image could be:”. It is important to note that this structural pattern does not result from the fusion of the two captions. Rather, it is a deliberate prompt used by the Davinci model to provide context and generate relevant responses. The third aspect is not a true limitation of the proposed model but rather one of its unique characteristics. Some of the captions generated by the LLM model incorporate emotional elements, adding depth and richness to the captions. For example, phrases like “a trolley car driving down a city street at night creates a captivating scene” or “a green bench in a subway station with a yellow train in the background creates a vivid and vibrant scene”. These emotional aspects are not explicitly provided in the input captions but are instead generated by the LLM model.
@@ 7 Conclusion and Future work
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] We describe the contributions of the presented work in the abstract and at the end of Section 1.
    2. (b)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] Yes and we believe that our paper conforms to them.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] .
    4. (d)
      Did you describe the limitations of your work? [Yes] We dedicate a paragraph in Section 6 describing the limitations of our proposal.
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] All the necessary code is available at this link https://drive.google.com/drive/folders/1tXbJyvwrEF-a-t-NZbt2xjoKaiGGeaz1?usp=sharing
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Yes. For the experimental results of finetuning OFA across multiple data splits we reported the standard deviation.
    4. (d)
      Did you include the amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Appendix A.3.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We cite the articles in which the exploited methods were presented and in Appendix A.1 we report links and details of the methods.
    2. (b)
      Did you mention the license of the assets? [Yes] In Appendix A.1 we mention the license of the used assets.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [No]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Section 6.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] We describe the purpose of the subjective study and provide a screenshot of the used interface in Section 4.2.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Cost and time estimation for subjective studies are discussed in Appendix B.4.
Appendix A Implementation Details
In this section we describe the selected models used for caption generation. Next we discuss the choices that led to the definition of the prompt for Davinci. Finally, the resources used for the development of the proposed work are detailed.
A.1 Captioning models
BLIP-2 Li et al. (2023) is a vision-language model that leverages pretrained Visual Language Processing (VLP) and NLP models to effectively handle tasks. It employs a Querying Transformer (Q-Former) to connect the image encoder and Large Language Model (LLM) and extract relevant features based on queries. The model undergoes two pretraining phases: Vision-and-Language Representation Learning and Vision-to-Language Generative Learning. The former aligns image and text representations through contrastive learning, grounded text generation, and matching, while the latter connects the Q-Former and frozen LLM. BLIP-2 achieved its best performance using ViT-g/14 as the Image Encoder and OPT2.7B as the LLM for image captioning on the MS-COCO dataset. ExpansionNet v2 Hu et al. (2022a) is an image captioning model that leverages expansion methods for improved performance. The model architecture is based on an encoder-decoder architecture built on top of a Swin Transformer in the Large configuration (SwinT-L) Liu et al. (2021). In this model, the image undergoes encoding using SwinT, followed by Block Static Expansion. During training, the ground-truth caption is encoded using a FeedForwardNet, which incorporates a Dynamic Expansion module. Visual and text representations are then combined using Multi-Head Attention and further processed by a classifier. The model parameters are efficiently learned using an end-to-end alternating full and partial training strategy. GIT Wang et al. (2022a) is a vision-language model that combines an image encoder and a text decoder. The image encoder is based on a Swin-like vision transformer pretrained on image-to-text data. It produces normalized and projected features, serving as input to the text decoder. The text decoder, a Transformer network, generates text descriptions using self-attention and feed-forward layers. The input text is processed and combined with visual features for decoding. Pretraining involves mapping images to their associated text descriptions, using a language modeling objective. The model was trained on 800M image-text pairs, employing a language modeling task for both pretraining and finetuning. OFA Wang et al. (2022b) is a versatile model designed for various vision-language tasks. It follows a task- and modality-agnostic approach, supporting both unimodal and multi-modal tasks. Image and text representations are unified in a token-based space, including object tracking and coordinates. Specifically, an image is encoded with ResNet modules to obtain patch-level embeddings, a text sequence is first tokenized using Byte-Pair Encoding (BPE) Sennrich et al. (2016) then subword tokens are encoded to embeddings. The encoder-decoder architecture consisting of Transformers is pretrained on vision-language, vision, and language datasets using cross-entropy loss. For image captioning, two finetuning steps optimize cross-entropy loss and CIDEr metric. ViT-GPT2 NLP Connect (2022) is an image captioner consisting of a ViT as the image encoder and the GPT2 as the text decoder. ViT processes images by dividing them into patches using attention mechanisms, while GPT2 analyzes and generates natural language text. Both models are pretrained, making implementation straightforward and achieving good performance. The ViT-GPT2 model combines the strengths of these unimodal models, providing a simple yet effective multi-modal approach for image captioning tasks.
The source code for BLIP-$2$https://github.com/salesforce/LAVIS (last access: 14 June 2023) is released under the BSD-3-Clause license, the source codes for ExpansionNet-v2333https://github.com/jchenghu/ExpansionNet_v2.git (last access: 14 June 2023), OFA [^4] https://github.com/OFA-Sys/OFA.git (last access: 14 June 2023), GIT [^5] https://github.com/huggingface/transformers.git (last access: 14 June 2023), and ViT-GPT2666https://huggingface.co/nlpconnect/vit-gpt2-image-captioning/tree/main (last access: 14 June 2023) are available under the Apache-2.0 license. The MS-COCO dataset is distributed under a CC-BY 4.0 license.
A.2 The Davinci model
The Davinci model is integrated into our captioning model using the OpenAI API [^7] https://openai.com/blog/openai-api. Interaction with the API is done through HTTP requests via official Python bindings. The OpenAI API requires API keys for authentication [^8] https://platform.openai.com/account/api-keys.
  • Explicitly instructing the model to fuse sentences both syntactically and semantically facilitates the creation of a result that goes beyond a mere summary of the two captions. Instead, it produces a genuine amalgamation of information while eliminating redundancy in terms and/or content. For instance, if both analyzed sentences feature the term “kite” in a specific syntactic role, the model comprehends that there is a single subject and avoids duplicating the term “kite” when generating the final caption.
A.3 Resource used
To perform the experiments for this work, namely running SoTA captioners and finetuning the OFA model, we exploit a desktop computer with an Intel Core i7-7700 CPU@3.60GHz, 16 GB DDR4 RAM 2400 MHz, and NVIDIA Titan X Pascal with 3840 CUDA cores. The OpenAI API is exploited for the caption fusion step of the proposed model.
Appendix B Analysis
In this section, we begin by conducting an analysis of BLIPScore, followed by presenting qualitative results to highlight the potential of the proposed captioning model. As part of an ablation study, we evaluate the performance of the proposed model using the base version, wherever available for SoTA models. Lastly, we delve into the configuration and outcomes of the study conducted with AMT.
B.1 BLIPScore analysis
In this analysis, we examine the distribution of BLIPScores across the 5000 test images from MS-COCO. Figure 6 displays several boxes representing different captioning models. These include five SoTA models, captions generated by our proposed model, captions with the highest BLIPScore among the five SoTA captioners, and human-made captions. The plot reveals several noteworthy findings. First, our proposed model achieves a slightly higher median score compared to the “Our (Best)” solution, indicating that our model incorporates additional details that enhance the alignment between the description and the image. Second, the ViT-GPT2 model performs the poorest, displaying the least accurate representation of image content. Lastly, it is interesting to observe the high interquartile range of BLIPScores for ground-truth captions. This characteristic arises due to the substantial variability in the estimated BLIPScores. Further examination of images with low BLIPScores revealed that some user-written captions lack detail and are inadequate.
Figure 6: Boxplot presenting the BLIPScore distribution for the 5000 test images with respect to generated and ground-truth captions.
In Figure 7, we present the percentage of SoTA model pairs selected during the ranking phase. In practice, these results demonstrate which models most frequently belong to the captions that are fused by our model. As shown, the pair of models BLIP-2 and OFA is selected for 22% of test image captions, while GIT and OFA are selected for 19%. The pair consisting of ExpNet-v2 and ViT-GPT2 accounts for only 2.4% of the generated captions.
B.2 Qualitative results
Figure 8 showcases images accompanied by their respective ground-truth and generated captions. From the provided examples, it is evident that our model excels in producing more detailed captions. For instance, in the image featuring a child with a toothbrush, our model seamlessly integrates information about the toothbrush, overalls, and hands. Similarly, in the image of the tower with a clock, our model enriches the description with intriguing qualitative details such as the presence of red and gold. On the other hand, as consistent with the discussions presented in the section outlining the limitations of our model, when the captions generated by the SoTA models exhibit significant similarity, our generated captions may not provide additional details (as observed in the image featuring the pole on the sidewalk). In some cases, the generated caption might also include redundant information, as seen in the description of the ducks in the pond, where terms like “muddy pond” and “brown water” are used together.
Refer to caption Ground-truth captions: A street sign for queen street with a colorful representation of the street on the pole. A decorated street sign for Queen Street with two blue signs below. A close up a street pole with a homemade street sign. We are looking at a street sign at the corner. A pole holding the street sign for Queen Street is decorated with a painting of a queen. Captions generated by existing models: BLIP-2: A street sign on a pole on the side of the road. ExpNet-v2: A wooden pole with street signs on the side of it. GIT: A street sign on a pole with a blue sign on it. OFA: A street sign on a pole on the side of a street. ViT-GPT2: A street sign on a pole on a sidewalk. Caption generated by the proposed model: Our: A wooden pole with street signs mounted on the sidewalk.
Refer to caption Ground-truth captions: A infant holding a baby toothbrush in his hand looking at it. A small baby is holding a white and blue toothbrush. A child holds a toothbrush in their hand. The baby is holding and looking at his tooth brush. A baby holds a toothbrush in its hand. Captions generated by existing models: BLIP-2: A baby in overalls holding a toothbrush. ExpNet-v2: A baby holding a toothbrush in his hand. GIT: A baby holding a toothbrush in his hands. OFA: A baby holding a toothbrush in its hands. ViT-GPT2: A baby holding a toothbrush in its mouth. Caption generated by the proposed model: Our: A baby in overalls is holding a toothbrush in his hands.
Refer to caption Ground-truth captions: Two Mallards float and drink water in a shallow pool. Two ducks in a pond in a camp grounds. Two ducks swimming in a body of water. Two ducks swimming on a very dirty pond. Two ducks are swimming in some brown water. Captions generated by existing models: BLIP-2: Two ducks swimming in a pool of brown water. ExpNet-v2: Two ducks swimming in the water in a pond. GIT: Two ducks swimming in a muddy pond with a car in the background. OFA: A couple of ducks swimming in a body of water. ViT-GPT2: A duck swimming in a pond with a duck in it. Caption generated by the proposed model: Our: Two ducks swimming in a muddy pond with a car in the background, amidst a pool of brown water.
Refer to caption Ground-truth captions: A large tall tower with a clock on the top. A red, white and gold decorative pillar with a clock on top. Pedestrians rest at the base of a large, public ornamental clock. A large tower in a crowded area with a clock at the top. A tall red and white clock tower in the middle of a street. Captions generated by existing models: BLIP-2: A red and gold tower with a clock on top. ExpNet-v2: A red clock tower in the middle of a street. GIT: A large clock tower with a large clock on it. OFA: A red clock tower with people sitting around it. ViT-GPT2: A clock tower with a clock on it’s side. Caption generated by the proposed model: Our: A red and gold clock tower with people sitting around it.
B.3 Evaluation of base model versions
The captioning pipeline designed was also implemented using the reduced versions of the SoTA image captioning models, i.e. the Base versions of GIT and OFA. As done for the more expensive model versions, also in this case the evaluation was conducted on the MS-COCO test set. Table 5 contains the quality and diversity results of the generated captions. As expected, the performance achieved using the reduced versions of the SoTA models is lower compared to the results reported in Table 2. This difference is particularly noticeable in the captioning metrics. However, it is worth noting that the BLIPScore remains almost identical, with a value of 0.65 for OFAHuge and 0.62 for OFABase. The BLIPScore obtained by our model with the best SoTA versions is 0.70 and 0.69 with the Base versions. These findings suggest that in terms of caption adequacy, the variance between the two model versions has minimal impact.
Table 5: Captioning performance by involving the Base version of GIT and OFA. Performance is reported in terms of image-agnostic metrics together with diversity statistics on the Karpathy split of MS-COCO. For each metric, best and second best methods are highlighted in bold and underlined, respectively.
Model Quality Diversity
B@4 ($\uparrow$) M ($\uparrow$) C ($\uparrow$) S ($\uparrow$) BLIPScore ($\uparrow$) mBLEU ($\downarrow$) Div-1 ($\uparrow$) Div-2 ($\uparrow$)
BLIP 39.8 31.0 133.4 23.8 0.62 0.78 0.47 0.30
ExpNet-v2 40.7 30.0 139.5 24.4 0.59 0.73 0.44 0.28
GITBase 36.7 28.5 122.9 22.0 0.59 0.70 0.43 0.28
OFABase 42.6 31.3 146.4 25.5 0.62 0.80 0.46 0.30
ViT-GPT2 35.4 27.9 119.1 21.2 0.48 0.59 0.42 0.27
Our (Best) 38.3 30.6 134.3 24.6 0.68 1.00 0.47 0.30
Our (Fusion) 28.0 30.9 98.5 25.6 0.69 0.50 0.60 0.40
B.4 AMT analysis
We had no experience with AMT prior to this study. Initially, we enlisted generic workers to rate each of the 1000 images, compensating them with \$0.02 per evaluation. The remuneration was determined based on the assumption that each assessment would require a maximum of 90 seconds. In total, we allocated \$90 for collecting assessments from the generic workers. We later acquired the judgments from experienced workers who we paid \$0.10 for a total of \$330.
Figure 9 illustrates the level of agreement among the six workers involved in the subjective study when selecting the best caption for each image. The graph addresses the question of how many of the 1000 images received a unanimous consensus among the workers. As observed from the results, there is a general consensus among the workers that our captions are superior. Specifically, for 308 images, two workers agreed on our caption as the best, for 266 images, three workers agreed, for 118 images, four workers agreed, finally, for 31 images, five workers chose our caption.

### v3

@@ 5 Evaluation
In this section, we investigate whether the best SoTA model, specifically OFA, can learn to generate more descriptive captions when trained using our generated captions as ground-truths. To conduct this experiment, we finetune the model for the image captioning task on MS-COCO by replacing the original ground-truth captions with the captions generated by our model. Due to limited computational resources – a desktop computer with an Intel Core i7-7700 CPU@3.60GHz, 16 GB DDR4 RAM 2400 MHz, and NVIDIA Titan X Pascal with 3840 CUDA cores – we utilize the Base version of OFA and adhere to the setup parameters recommended by the authors. The 5000 samples from the MS-COCO test set are randomly divided into training (3500 samples), validation (500 samples), and testing (1000 samples). To minimize the bias introduced by partitioning, we repeat this split process five times. Table 8 reports the quality captioning performance of OFABase trained on the original MS-COCO and finetuned on MS-COCO with our ground-truths, respectively. It is possible to see that the results for the finetuned model are slightly higher than the baseline ones, i.e. 2 for B@4, 1 for M, and 10 for C, respectively.
Table 8: Captioning performance comparison between two versions of OFABase, i.e. the model pretrained on MS-COCO (also known as the baseline) and the model finetuned on the dataset with our generated captions used as ground-truth. It is important to note that the reported performances represent the mean and standard deviation across 5 train-val-test splits.
@@ 6 Discussion
Existing datasets and caption models have several limitations, such as short ground-truth captions and generated captions that tend to look alike and are semantically poor. In this article, we propose a caption model that tries to overcome the previous limitations by generating longer and more detailed captions without requiring human intervention. The captions generated by our model were evaluated from different points of view. Here, we summarize the contributions of our work, answering the research questions posed in the Introduction. In Section 4.2, we address [RQ1] and demonstrate that the captions generated by our proposed model outperform those generated by the SoTA models in terms of accuracy, diversity, and semantic richness. In Section 4.3, we tackle [RQ2] through a subjective study involving human participants. Based on human judgment, our model generates captions that are more appealing than those produced by SoTA models. In Section 5, we aim to answer [RQ3]. The results of our experiment highlight how current image captioning models lean towards the use of common words, phrases, and linguistic patterns. Consequently, this leads to a dearth diversity in the generated captions, even when the models are trained on more elaborate and descriptive caption datasets [8]. This phenomenon, known as the mode collapse problem, has been widely discussed in generative modeling [53]. Addressing this issue requires further investment. Consequently, our proposed caption model can serve as a valuable resource for generating ground-truth captions and pushing models beyond their current limitations.
6.1 Societal impact
@@
Compared to previous image captioning models, our model demonstrates improvements in performance and is more suitable for helping people with visual impairments. The models that are exploited by our proposal are pretrained on large-scale data, and the data are not guaranteed to contain no toxic content, which could negatively impact the output generated. While we have observed only a small number of such cases in our qualitative analysis, it is imperative to exercise caution when deploying the model into practical applications. In addition, more research investigations are needed to ensure better control over model output.
6.2 Limitations
@@
The limitations of the proposed model are mainly three. The first is that sometimes the five SoTA captioners generate identical or very similar captions (this occurs for about 150 of the 5000 images in the MS-COCO test set). In this case, the fusion of the captions results in a caption lacking additional detail. An example is the following where BLIP-2, ExpNet-v2 and OFA generate the same caption, namely “a giraffe standing in the middle of a field”, GIT describes the image content as “a giraffe standing in a field with trees in the background”, finally ViT-GPT2 predicts that “a giraffe standing in a field with trees”. The caption generated by our model is “a giraffe stands majestically in the middle of a field”. Secondly, for 16 test images, the LLM model generates captions with the prefix “The caption for the image could be:”. It is important to note that this structural pattern does not result from the fusion of the two captions. Rather, it is a deliberate prompt used by the Davinci model to provide context and generate relevant responses. The third aspect is not a true limitation of the proposed model but rather one of its unique characteristics. Some of the captions generated by the LLM model incorporate emotional elements, adding depth and richness to the captions. For example, phrases like “a trolley car driving down a city street at night creates a captivating scene” or “a green bench in a subway station with a yellow train in the background creates a vivid and vibrant scene”. These emotional aspects are not explicitly provided in the input captions but are instead generated by the LLM model.
@@ 7 Conclusion and Future work
Declarations
Not applicable.